OpenStudy (anonymous):

Please Help! WILL GIVE MEDAL!! Cameron is replacing the roof on a gazebo. The roof is in the shape of a cone with an open bottom. The radius is 7 feet and the slant height is twelve feet. What is the lateral area of the roof? Use three point one four for pi. two hundred sixty-three point seven six square feet three hundred seventy-six point eight square feet seven hundred fifteen point nine two square feet eight hundred seventy-nine point two square feet

OpenStudy (anonymous):

lateral area = pi*r*s where s is the length of the slope of the cone pi*r*s = pi*10*6 = 60*pi = 188.4956 ft^2 this is what i found .

OpenStudy (anonymous):

That's not an option though:(

OpenStudy (anonymous):

Ok......so The formula is πrh. Where: π = 3.14 r = Radius h = Height...Hope you can take it from here?

OpenStudy (anonymous):

so 3.14 times 7 times 12= 263.76... thanks!!!!

OpenStudy (anonymous):

Let me write for you anyways: πrh (3.14)(7)(10) (It gives you the diameter, and the radius is half the diameter which gives us 7) (21.98)(10) 219.8